OpenStudy (anonymous):

Solve for the variable x: x + 3 = 6 x + 2 5

OpenStudy (saifoo.khan):

is the ques correct?

OpenStudy (anonymous):

its messed up. h/on

OpenStudy (anonymous):

x + 3 = 6 x = 6-3 x=3 now x+25 = 3+25 = 28

OpenStudy (saifoo.khan):

x + 3 ------ = 6 x + 25 this?

OpenStudy (anonymous):

its 6 over 5 so the bottom would be :x +2

OpenStudy (saifoo.khan):

x + 3 6 ------ = ----- x + 2 5

OpenStudy (anonymous):

yes

OpenStudy (saifoo.khan):

cross multiply. 5(x + 3) = 6(x + 2) 5x + 15 = 6x + 12 Shuffle: x = 15 -12 x = 3